Find the probability of drawing a heart or a face card from a standard deck of cards.: Find the probability of drawing a heart or a face card from a standard deck of cards.
Answer by Fombitz(1740)

52 cards in a deck
K,Q,J face cards in a deck - 12 total
P(face)=12/52=3/13
.
.
.
13 hearts in a deck
P(heart)=13/52
.
.
.
P(face card and heart) = 3/32
.
.
.
Probability of either a face card or a heart, add individual probabilities, but remember to subtract the three heart face cards, to not count them twice.
P(A or B)=P(A)+P(B)-P(A and B)
where A is drawing a face card and B is drawing a heart.
.
.
.
P(face or heart)=P(face)+P(heart)-P(face and heart)
P(face or heart)=12/52+13/52-3/52=22/52=11/26